OpenStudy (anonymous):

Solve,The speed of a car uphill is 56 mph, speed downhill 72 mph, speed in level ground 63 mph. The time taken to travel from Town A to Town B is 4hrs, and back from Town B to Town A is 4 hours 40 minutes. What is the distance between A and B?

OpenStudy (anonymous):

Heading out, x/72 + y/63 + z/56 = 4 Returning, x/56 + y/63 + z/72 = 14/3 Now yu multiply equations(56, 63, and 72) by the LCM of denominators 7x+8y+9z= 4*7*8*9 9x+8y+7z=(14/3)*7*8*9 So now, 16(x+y+z)=(26/3)*7*8*9 & what does that equal.?

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Did yu get the answer.?

OpenStudy (anonymous):

ya,i got those 2 equations.i was looking for third equation to fing x,y and z separately.Now i found that it can be solved directly.

OpenStudy (anonymous):

Alright good luck with the rest.! ^^